# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v1

@@ Abstract
Do home prices incorporate flood risk in the immediate aftermath of specific flood events, or is it the repeated exposure over the years that plays a more significant role? We address this question through the first systematic study of the Italian housing market, which is an ideal case study because it is highly exposed to floods, though unevenly distributed across the national territory. Using a novel dataset containing about 550,000 mortgage-financed transactions between 2016 and 2024, as well as hedonic regressions and a difference-in-difference design, we find that: (i) specific floods do not decrease home prices in areas at risk; (ii) the repeated exposure to floods in flood-prone areas leads to a price decline, up to 4% in the most frequently flooded regions; (iii) responses are heterogeneous by buyers’ income and age. Young buyers (with limited exposure to prior floods) do not obtain any price reduction for settling in risky areas, while experienced buyers do. At the same time, buyers who settle in risky areas have lower incomes than buyers in safe areas in the most affected regions. Our results emphasize the importance of cultural and institutional factors in understanding how flood risk affects the housing market and socioeconomic outcomes.
22footnotetext: We would like to thank Paolo Bajardi, Diana Bruni, Matteo Coronese, Paola Danelutti, Adriano di Fabio, Barbara Lenti, Jacopo Lenti, Laura Li Puma, Michela Manuelli, Carla Monferrato, Yamir Moreno, Silvia Ronchiadin, Simone Scarsi, and Fabio Verachi for their helpful comments at various stages of this work.11footnotetext: Equal contribution.33footnotetext: Corresponding author: marco.pangallo@centai.eu.
@@ 1 Introduction
Research on the impact of flood risk on home prices, primarily based on U.S. data, typically finds small effects, except in the years immediately following major flood events when the price impact in affected areas is more pronounced. For example, prior to Hurricane Sandy, flood risk had little influence on home prices in New York City. However, in the years after the storm, properties in flood zones –despite being undamaged– sold for up to 8% less than properties in safer areas (Ortega and Taṣpınar, 2018). Bin and Landry (2013) attribute this pattern to the availability heuristic (Tversky and Kahneman, 1973), suggesting that individuals must experience floods firsthand to perceive them as a significant threat and demand a price discount for homes at risk.
@@
In this paper, we propose an alternative yet complementary hypothesis, which we term historical memory. We suggest that, in places with storms not as destructive as hurricanes and different housing market institutions than in the U.S., isolated flood events may be insufficient to significantly shift market beliefs and prompt the pricing of flood risk. Instead, repeated flood exposure in the same region may have a more substantial influence. If this hypothesis is true, home prices in frequently flooded areas are likely to reflect flood risk more prominently than those in rarely flooded regions, even for properties that remain undamaged. Moreover, younger buyers, having experienced fewer floods, may require a smaller discount to purchase at-risk properties compared to more experienced buyers.
@@
We test this hypothesis through the first systematic study of the Italian housing market, which offers an ideal case study for several reasons. First, Italy faces significant flood risk, with over 10% of its population and buildings exposed to floods with a return period of less than 200 years (Trigila et al., 2021), and approximately 40 major floods recorded since 2000 (EM-DAT, 2024). Second, Italy recently experienced a major flood in the Emilia-Romagna region on May 16–17, 2023, which displaced 45,000 people and caused damages estimated at around 10 billion euros (EM-DAT, 2024). Third, flood exposure varies across regions: while flood risk affects the entire country, Northern Italy experiences more frequent events than the South. Fourth, unlike markets in the U.S. or the U.K., Italy lacks institutional requirements for flood risk disclosure during property sales and neither mandates nor promotes flood insurance. Finally, floods in Italy typically result from storms with a less catastrophic impact than the hurricanes or tropical storms common in the U.S.
@@
Given this environmental and institutional context, we frame and test our historical memory hypothesis as follows. We argue that, without an institutional structure that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. Due to the relatively limited destructive potential of individual floods, a single event may not significantly shift market beliefs. However, repeated flood occurrences gradually reinforce risk awareness, eventually leading to its capitalization in home prices. The recent Emilia-Romagna flood provides an opportunity to test the first aspect of this hypothesis, while the variation in flood risk across Italian regions enables us to examine the second.
@@
The main reason why there is so little research on how flood risk impacts the Italian housing market is that systematic data only started to exist in the last few years. [^1] We are only aware of one study assessing the impact of flood risk on prices (Rosato et al., 2017), using 197 data points manually collected from real estate agencies in Venice. Only recently systematic data about the housing market became available from the housing advertisement website Immobiliare.it (Pangallo and Loberto, 2018; Loberto et al., 2022), but concerning our research question, these data have only been used to study the impact of a sea wall in Venice (Benetton et al., 2023). In this paper, we take advantage of a novel database that contains the near-universe of mortgages issued by Intesa Sanpaolo, the largest Italian bank. Our final dataset contains about 550,000 mortgages issued from 2016 to 2024, which is about 15% of Italian mortgage-financed transactions, and 10% of total transactions. [^2] The data are highly representative of the Italian housing market throughout the entire national territory, see Supplementary Section A.2. For each transaction, we know the sale price, the date, the income of the buyer, whether the buyer is under-36, and several physical characteristics of the transacted home, including its geographical coordinates. We use the coordinates to determine whether a home is inside a flood-prone area by performing spatial matching with risk maps provided by ISPRA, the Italian Institute for Environmental Protection and Research (Trigila et al., 2021). Using the same approach, we also assess whether homes have been hit by specific flood events, such as the 2023 Emilia-Romagna flood. By “hit”, we mean that their coordinates lie within the maximal flooding area identified by the Copernicus EU program (Directorate Space, Security and Migration, European Commission Joint Research Centre , EC JRC).
@@
First, we assess whether specific flood events lead to stronger price penalties in the affected areas. Most studies find price reductions in a 5-25% range, that only last a few years after the event, and interpret these findings as evidence for a transitory increase in attention to flood risk. [^3] See Hallstrom and Smith (2005); Daniel et al. (2009b); Kousky (2010); Atreya et al. (2013); Bin and Landry (2013); Atreya and Ferreira (2015); Ortega and Taṣpınar (2018); Beltrán et al. (2019). In line with most of the literature, we follow a difference-in-difference approach that distinguishes between homes at risk and homes not at risk, as well as homes sold before or after the event. In addition to most of the literature, we further distinguish between homes that were hit by the flood and homes that were not hit but could have been, as they were in risky areas in the same municipalities as hit homes (Ortega and Taṣpınar, 2018). We find that the 2023 Emilia-Romagna flood (which, as discussed, has been the largest flood in Italy in the last 20 years) did not lead to any further price reduction for homes at risk that were not hit. [^4] We also find that prices of hit homes decreased by 10% in the first 9 months after the event, reversing to a 10% appreciation between 9 and 15 months after the event. Most likely, affected homes were first damaged and then renovated, resulting in a price premium after renovations ended. This result, which contrasts with much of the literature, is robust across specifications, home types (e.g., focusing on ground floor apartments), and considering the second and third largest flood events in the studied period.
@@
Second, we assess whether flood risk reduces prices, independently of specific flood events, and how this varies across regions with different levels of exposure. All else equal, economic theory predicts a price penalty for homes at risk (Bin and Landry, 2013). However, because of confounders, such as water-related amenities (Bin and Kruse, 2006; Daniel et al., 2009a), it is hard to establish causality in empirical studies, [^5] For instance, nice views over rivers or lakes may both increase prices and be correlated to high flood risk. especially given that we cannot rely on repeated sales (Beltrán et al., 2019) or changes in risk maps (Hino and Burke, 2021). [^6] In this work, we adopted the 2020 release of the ISPRA risk maps. A 2017 release is available but does not present major changes in the total risk areas with respect to the 2020 version. To address this issue, we run hedonic regressions that take into account multiple physical characteristics of transacted homes, as well as granular spatial fixed effects, down to the census tract level. We argue that these fixed effects capture enough spatial variation so that the coefficient on flood risk is likely to capture the actual penalty of being located in a flood zone. We find that homes at risk sell at approximately 1% less than homes that are not at risk. This risk penalty is not at all uniform across the Italian territory. It is only statistically significant in the most frequently flooded regions, reaching 2.5% in Liguria, 3.5% in Emilia-Romagna and 4% in Toscana.
@@
Third, we study how these results vary across buyers with different age and income. We find that young buyers [^7] More specifically, under-36 buyers with an adjusted income (“ISEE”) less than 40,000€  receiving a mutuo giovani, namely a mortgage with favorable conditions guaranteed by the government. do not obtain any discount for purchasing homes at risk, while more experienced buyers obtain a discount that increases with the frequency of floods in their region. High income buyers, whether young or not, are also more likely to obtain a discount. Furthermore, in line with Zivin et al. (2023), we find that lower income buyers are more likely to purchase properties in risky areas, in part because of lower prices. As before, this effect is statistically significant only in the most frequently flooded regions.
@@
Summing up, in the immediate aftermath of the largest Italian floods in the last 20 years, we do not observe any price reductions for homes that were in the affected municipalities even though not directly hit. At the same time, we find a small “baseline” penalty of flood risk on prices, but this penalty increases up to four times in the most flood-prone parts of Italy. We interpret this finding as evidence for our hypothesis of “historical memory”, whereby it is not isolated floods, but the repeated exposure to floods, that drives the incorporation of flood risk in home prices. This evidence is further corroborated by our finding that only experienced buyers, who may have witnessed multiple floods in their lifetime, obtain a discount for purchasing homes at risk. Finally, our findings on the income of buyers moving to at-risk areas have welfare implications, and can be used to design effective policies.
@@
A first question is how much specific flood events heighten housing market participants’ perceptions about climate risk, leading to stronger price penalties for homes located in risk zones. Lots of studies address this question (Hallstrom and Smith, 2005; Daniel et al., 2009b; Kousky, 2010; Atreya et al., 2013; Bin and Landry, 2013; Atreya and Ferreira, 2015; Ortega and Taṣpınar, 2018; Beltrán et al., 2019), generally finding a temporary increase in the price penalty that only lasts few years. [^8] Zivin et al. (2023) are among the few that find a price premium in the years following hurricanes, suggesting that it may be due to a supply shortage due to destroyed homes. For instance, Bin and Landry (2013) argue that the availability heuristic (Tversky and Kahneman, 1973) explains why individuals need to experience flood events to attach a non-negligible probability to their occurrence. They find that two hurricanes occurring in 1996 and 1999 in North Carolina, after no major hurricane since 1954, lead to price reductions of 5.7% and 8.8%, respectively. Atreya and Ferreira (2015) argue that it is important to distinguish between inundated and non-inundated homes, because price penalties may conflate the physical damage that occurred to homes and the increased attention that buyers may pay to flood risk. Ortega and Taṣpınar (2018) are able to clearly distinguish between the two cases, showing that price penalties in New York homes damaged by Hurricane Sandy were up to 17-22% for several years, while penalties for homes at risk that were not damaged were around 8%. Our results also show a strong decrease in prices of affected homes, around 10% between 3 and 9 months after the Emilia-Romagna flood, but reconstruction is much faster, as prices appear to recover already after 9 months (presumably, this is linked to the difference in intensity between Hurricane Sandy and the Emilia-Romagna floods). Differently from most of the literature, however, we do not find any significant change in the price penalty for homes that were not directly hit by the flood, but still located in risky areas of affected municipalities.
@@
The next question is how much flood risk is capitalized in home prices in “normal times”, i.e. not in the immediate aftermath of specific flood events. [^9] Several studies use spatial regressions to estimate these effects (Atreya et al., 2013; Bin and Landry, 2013; Zhang, 2016; Pommeranz and Steininger, 2020). Here, we do not find relevant spatial correlations after controlling for geographical units, as evidenced by Global Moran’s I tests and LISA cluster maps, so we did not follow this direction. See Supplementary Section B.2. Almost all previous studies relied on U.S. data. Daniel et al. (2009a) review estimates that range from a -52% penalty for at-risk homes to a +58% premium (mostly due to failing to control for water-related amenities), with a mean estimate of -2.3%. Beltrán et al. (2018a) update these estimates considering more studies, and find a range between a -75.5% penalty and 61% premium, with an adjusted mean -4.6% penalty for inland flooding. More recently, in what is perhaps the most complete study on this question, Hino and Burke (2021) find smaller -2.1% and -1.4% penalties with panel data and diff-in-diff approaches, respectively. Using actuarially fair insurance costs, they argue that these discounts are much smaller than the ones that would be expected in efficient markets, ranging from -4.7% and -10.6%. Our average 1% price penalty for homes located in flood zones relates to these findings. Unsurprisingly, we find a smaller penalty compared to the U.S., as in Italy disclosure of flood risk is non-existent and flood insurance is in many cases not even encouraged. [^10] It is harder to say to what extent flood risk is capitalized in home prices, because although the penalty is smaller in Italy than in the U.S., the extent of damages is also likely to be smaller, as Italian storms are less strong than hurricanes in the U.S.
@@
The third piece of literature that we relate to is the one looking at the socio-economic characteristics of the buyers. Although there is substantial evidence that “sophisticated” buyers, such as non-owner-occupiers (Bernstein et al., 2019), or “believers” in climate change (Baldauf et al., 2020), obtain higher discounts when purchasing at-risk properties, the evidence on the effect of age or income is more scarce. For instance, Bradt et al. (2021) find that younger and lower-income buyers are less likely to purchase flood insurance, but Fairweather et al. (2024) do not find any significant difference in terms of searching at-risk properties across income levels. We contribute to this literature by showing that, in our case, younger and lower-income buyers obtain a smaller discount than more experienced and higher-income buyers. Regarding the income of buyers who settle in at-risk areas, our main comparison is Zivin et al. (2023), who match information on buyers’ income with transactions. They show that periods in which prices increase correspond to periods in which higher-income buyers are more likely to purchase properties in affected areas. We find a similar result concerning incomes, as we find that areas at risk show slightly lower home prices, thus households with slightly lower income move there. Ellen and Meltzer (2024) also find that lower-income households moved into the areas of New York that were hit harder by Hurricane Sandy.
@@
First, Section 2 describes the data from Intesa Sanpaolo, the data from ISPRA and Copernicus spatially matched to transacted homes, as well as EM-DAT dataset used to define the regional frequency of flood events. Next, Section 3 describes our econometric approach. The main results are in Section 4. Section 5 concludes.

### v2

@@ Abstract
Do home prices incorporate flood risk in the immediate aftermath of specific flood events, or is it the repeated exposure over the years that plays a more significant role? We address this question through the first systematic study of the Italian housing market, which is an ideal case study because it is highly exposed to floods, though unevenly distributed across the national territory. Using a novel dataset containing about 550,000 mortgage-financed transactions between 2016 and 2024, as well as hedonic regressions and a difference-in-difference design, we find that: (i) specific floods do not significantly decrease home prices in areas at risk; (ii) the repeated exposure to floods in flood-prone areas leads to a price reduction, up to 4% in the most frequently flooded regions; (iii) young, low-income buyers are more likely to settle in flood-prone areas, but do not obtain a discount, while more experienced buyers do.
22footnotetext: We would like to thank Paolo Bajardi, Diana Bruni, Matteo Coronese, Paola Danelutti, Adriano di Fabio, Barbara Lenti, Jacopo Lenti, Laura Li Puma, Michela Manuelli, Carla Monferrato, Yamir Moreno, Silvia Ronchiadin, Simone Scarsi, and Fabio Verachi for their helpful comments at various stages of this work, as well as seminar participants at EAERE and IAERE, where this work was selected among the top 5 presentations by young researchers, to be included in a special session at the Italian Economic Society 66th Annual Conference.11footnotetext: Equal contribution.33footnotetext: Corresponding author: marco.pangallo@centai.eu.
JEL Codes: D83, Q51, Q54, R31
Keywords: Flood risk; housing markets; hedonic regressions; awareness
@@ 1 Introduction
Research on the impact of flood risk on home prices, primarily based on U.S. data, typically finds small impacts, except in the years immediately following major flood events, when prices substantially decline in affected areas. For example, prior to Hurricane Sandy, flood risk had little influence on home prices in New York City. However, in the years after the storm, properties in flood zones –despite being undamaged– sold for up to 8% less than properties in safer areas (Ortega and Taṣpınar, 2018). Bin and Landry (2013) attribute this pattern to the availability heuristic (Tversky and Kahneman, 1973), suggesting that individuals must experience floods firsthand to perceive them as a significant threat and demand a price discount for homes at risk.
@@
In this paper, we consider an alternative yet complementary hypothesis, which we term historical memory. We suggest that, in places with storms not as destructive as hurricanes and different housing market institutions than in the U.S., isolated flood events may be insufficient to significantly shift market beliefs and prompt the pricing of flood risk properties. Instead, repeated flood exposure in the same region may have a more substantial influence. If this hypothesis is true, home prices in frequently flooded areas are likely to reflect flood risk more prominently than those in rarely flooded areas, even for properties that remain undamaged. Moreover, experienced buyers, having witnessed more floods, may require a higher discount to purchase at-risk properties compared to younger buyers.
We operationalize the concept of historical memory by constructing a measure of buyers’ awareness of flood risk. This measure rises following each flood event but gradually declines as memories of past floods fade. A single flood, even if severe, may not raise awareness sufficiently for flood risk to become reflected in home prices, as buyers might perceive it as an isolated incident. In contrast, repeated exposure to flooding heightens awareness, making it more likely that flood risk is capitalized into property values. Our measure offers a concrete implementation of the ideas proposed by Tobin and Newton (1986) and Pryce et al. (2011), enabling empirical testing with real-world data.
@@
To test this hypothesis, we conduct the first systematic study of the Italian housing market, which is an ideal case study for several reasons. First, Italy faces significant flood risk, with over 10% of its population and buildings exposed to flood risk (Trigila et al., 2021), and approximately 40 major floods recorded since 2000 (EM-DAT, 2024). Second, Italy recently experienced a major flood in the Emilia-Romagna region on May 16–17, 2023, which displaced 45,000 people and caused damages estimated at around 10 billion euros (EM-DAT, 2024). Third, flood exposure varies across regions: while flood risk affects the entire country, Northern Italy experiences more frequent events than the South. Fourth, unlike markets in the U.S. or the U.K., Italy lacks institutional requirements for flood risk disclosure during property sales and neither mandates nor promotes flood insurance. Finally, floods in Italy typically result from storms with a less catastrophic impact than the hurricanes or tropical storms common in the U.S.
@@
Given this environmental and institutional context, we frame and test our historical memory hypothesis as follows. We argue that, without an institutional structure that encourages the sharing of information regarding flood risk, households rely on the historical memory of major floods to identify the most at-risk areas. Due to the relatively limited destructive potential of individual floods, a single event may not significantly shift market beliefs. However, repeated flood occurrences gradually reinforce risk awareness, eventually leading to its capitalization in home prices. The recent Emilia-Romagna flood provides an opportunity to test the first aspect of this hypothesis, while the variation in flood risk across the Italian territory enables us to examine the second.
@@
The main reason why there is so little research on how flood risk impacts the Italian housing market is that systematic data only started to exist in the last few years. [^1] We are only aware of one study assessing the impact of flood risk on prices (Rosato et al., 2017), using 197 data points manually collected from real estate agencies in Venice. Only recently systematic data about the housing market became available from the housing advertisement website Immobiliare.it (Pangallo and Loberto, 2018; Loberto et al., 2022), but concerning our research question, these data have only been used to study the impact of a sea wall in Venice (Benetton et al., 2023). In this paper, we take advantage of a novel database that contains the near-universe of mortgages issued by Intesa Sanpaolo, the largest Italian bank. Our final dataset contains about 550,000 mortgages issued from 2016 to 2024, which is about 15% of Italian mortgage-financed transactions, and 10% of total transactions. [^2] The data are highly representative of the Italian housing market throughout the entire national territory, see Supplementary Section A.2. For each transaction, we know the sale price, the date, the income of the buyer, a proxy for whether the buyer is under-36, and several physical characteristics of the transacted home, including its geographical coordinates. We use the coordinates to determine whether a home is inside a flood-prone area by performing spatial matching with risk maps provided by ISPRA, the Italian Institute for Environmental Protection and Research (Trigila et al., 2021). Using the same approach, we also assess whether homes have been hit by specific flood events, such as the 2023 Emilia-Romagna flood. By “hit”, we mean that their coordinates lie within the maximal flooding area identified by the Copernicus EU program (Directorate Space, Security and Migration, European Commission Joint Research Centre , EC JRC).
@@
First, we assess whether specific flood events lead to stronger price penalties in the affected areas. Previous studies, mostly based on the U.S., find price reductions in a 5-25% range, that only last a few years after the event, and interpret these findings as evidence for a transitory increase in flood risk awareness. [^3] See Hallstrom and Smith (2005); Daniel et al. (2009b); Kousky (2010); Atreya et al. (2013); Bin and Landry (2013); Atreya and Ferreira (2015); Ortega and Taṣpınar (2018); Beltrán et al. (2019). In line with most of the literature, we follow a difference-in-difference approach that distinguishes between homes at risk and homes not at risk, as well as homes sold before or after the event. In addition to most of the literature, we further distinguish between homes that were hit by the flood and homes that were not hit but could have been, as they were in risky areas in the same municipalities as hit homes (Ortega and Taṣpınar, 2018).
We find that the 2023 Emilia-Romagna flood (which, as discussed, has been the largest flood in Italy in the last 20 years) did not lead to any further price reduction for homes at risk that were not hit. [^4] We also find that prices of hit homes decreased by 10% in the first 9 months after the event, reversing to a 10% appreciation between 9 and 15 months after the event. Most likely, affected homes were first damaged and then renovated, resulting in a price premium after renovations ended. This result, which contrasts with much of the literature, is robust across specifications, home types (e.g., focusing on ground floor apartments), and considering the second and third largest flood events in the studied period.
@@
Second, we assess whether flood risk reduces prices, independently of specific flood events, and how this varies across parts of Italy with different levels of exposure. All else equal, economic theory predicts a price penalty for homes at risk (Bin and Landry, 2013). However, because of confounders, such as water-related amenities (Bin and Kruse, 2006; Daniel et al., 2009a), it is hard to establish causality in empirical studies, [^5] For instance, nice views over rivers or lakes may both increase prices and be correlated to high flood risk. especially given that we cannot rely on repeated sales (Beltrán et al., 2019) or changes in risk maps (Hino and Burke, 2021). To address this issue, we run hedonic regressions that take into account multiple physical characteristics of transacted homes, as well as granular spatial fixed effects, down to the census tract level. We argue that these fixed effects capture enough spatial variation so that the coefficient on flood risk is likely to capture the actual penalty of being located in a flood zone.
We find that homes at risk sell at approximately 1% less than homes that are not at risk. This risk penalty is not uniform. Dividing all transactions in our dataset by our measure of flood risk awareness in the area and at the time of the transaction, we find that flood risk is capitalized in home prices only in the top awareness tercile, where discounts reach 1.7%. The regions which averaged highest awareness over the considered period show stronger discounts, reaching 2.5% in Liguria, 3.5% in Emilia-Romagna and 4% in Toscana.
@@
Third, we study how these results vary across buyers with different age and income. We find that young buyers [^6] More specifically, under-36 buyers with an adjusted income (“ISEE”) less than 40,000€  receiving a mutuo giovani, namely a mortgage with favorable conditions guaranteed by the government. do not obtain any discount for purchasing homes at risk, while more experienced buyers obtain a discount that increases with our awareness measure. This result is an average across all income levels. Further disaggregating by income, we find that income plays a significant role only among young buyers, as high-income young buyers purchase homes at risk with a 3% discount compared to homes that are not at risk, while medium and low-income young buyers do not obtain any discount or even buy at a premium. Finally, in line with Zivin et al. (2023), we find that lower-income buyers are more likely to purchase properties in risky areas, in part because of lower prices.
@@
Summing up, in the immediate aftermath of the largest Italian floods in the last 20 years, we do not observe any price reductions for homes that were in the affected municipalities even though not directly hit. At the same time, we find a small “baseline” penalty of flood risk on prices, but this penalty increases up to four times in the most flood-prone parts of Italy. We interpret this finding as evidence for our hypothesis of “historical memory”, whereby it is not isolated floods, but the repeated exposure to floods, that drives the incorporation of flood risk in home prices. This evidence is further corroborated by our finding that only experienced buyers, who may have witnessed multiple floods in their lifetime, obtain a discount for purchasing homes at risk.
@@
A first question is how much specific flood events heighten housing market participants’ perceptions about climate risk, leading to stronger price penalties for homes located in risk zones. Lots of studies address this question (Hallstrom and Smith, 2005; Daniel et al., 2009b; Kousky, 2010; Atreya et al., 2013; Bin and Landry, 2013; Atreya and Ferreira, 2015; Ortega and Taṣpınar, 2018; Beltrán et al., 2019), generally finding a temporary increase in the price penalty that only lasts few years. [^7] Zivin et al. (2023) are among the few that find a price premium in the years following hurricanes, suggesting that it may be due to a supply shortage due to destroyed homes. For instance, Bin and Landry (2013) argue that the availability heuristic (Tversky and Kahneman, 1973) explains why individuals need to experience flood events to attach a non-negligible probability to their occurrence. They find that two hurricanes occurring in 1996 and 1999 in North Carolina, after no major hurricane since 1954, lead to price reductions of 5.7% and 8.8%, respectively. The fact that the second hurricane led to higher price reductions than the first supports our awareness measure. Atreya and Ferreira (2015) argue that it is important to distinguish between inundated and non-inundated homes, because price penalties may conflate the physical damage that occurred to homes and the increased attention that buyers may pay to flood risk. Ortega and Taṣpınar (2018) are able to clearly distinguish between the two cases, showing that price penalties in New York homes damaged by Hurricane Sandy were up to 17-22% for several years, while penalties for homes at risk that were not damaged were around 8%. Our results also show a strong decrease in prices of affected homes, around 10% between 3 and 9 months after the Emilia-Romagna flood, but reconstruction is much faster, as prices appear to recover already after 9 months (presumably, this is linked to the difference in intensity between Hurricane Sandy and the Emilia-Romagna floods). Differently from most of the literature, however, we do not find any significant change in the price penalty for homes that were not directly hit by the flood, but still located in risky areas of affected municipalities. [^8] An exception to the general findings of the literature is Kousky (2010), who do not find further reductions in 100-year floodplains in Missouri, as flood risk seemed to be already capitalized in home prices.
@@
The next question is how much flood risk is capitalized in home prices in “normal times”, i.e. not in the immediate aftermath of specific flood events. [^9] Several studies use spatial regressions to estimate these effects (Atreya et al., 2013; Bin and Landry, 2013; Zhang, 2016; Pommeranz and Steininger, 2020). Here, we do not find relevant spatial correlations after controlling for geographical units, as evidenced by Global Moran’s I tests and LISA cluster maps, so we did not follow this direction. See Supplementary Section B.2. Almost all previous studies relied on U.S. data. Daniel et al. (2009a) review estimates that range from a -52% penalty for at-risk homes to a +58% premium (mostly due to failing to control for water-related amenities), with a mean estimate of -2.3%. Beltrán et al. (2018a) update these estimates considering more studies, and find a range between a -75.5% penalty and 61% premium, with an adjusted mean -4.6% penalty for inland flooding. More recently, in what is perhaps the most complete study on this question, Hino and Burke (2021) find smaller -2.1% and -1.4% penalties with panel data and diff-in-diff approaches, respectively. Using actuarially fair insurance costs, they argue that these discounts are much smaller than the ones that would be expected in efficient markets, which would instead range from -4.7% to -10.6%. Our average 1% price penalty for homes located in flood zones relates to these findings. Unsurprisingly, we find a smaller penalty compared to the U.S., as in Italy disclosure of flood risk is non-existent and flood insurance is in many cases not even encouraged. [^10] It is harder to say to what extent flood risk is capitalized in home prices, because although the penalty is smaller in Italy than in the U.S., the extent of damages is also likely to be smaller, as Italian storms are less destructive than hurricanes in the U.S. As mentioned above, our estimates of price penalties depend on buyers’ awareness of flood risk, operationalizing ideas put forward by Tobin and Newton (1986) and Pryce et al. (2011).
@@
The third piece of literature that we relate to is the one looking at the socio-economic characteristics of the buyers. Although there is substantial evidence that “sophisticated” buyers, such as non-owner-occupiers (Bernstein et al., 2019), or “believers” in climate change (Baldauf et al., 2020), obtain higher discounts when purchasing at-risk properties, the evidence on the effect of age or income is more scarce. For instance, Bradt et al. (2021) find that younger and lower-income buyers are less likely to purchase flood insurance, but Fairweather et al. (2024) do not find any significant difference in terms of searching at-risk properties across income levels. We contribute to this literature by showing interactions between the age and income of the buyers. Regarding the income of buyers who settle in at-risk areas, our main comparison is Zivin et al. (2023), who match information on buyers’ income with transactions. They show that periods in which prices increase correspond to periods in which higher-income buyers are more likely to purchase properties in affected areas. We find a similar result concerning incomes, as we find that areas at risk show slightly lower home prices, thus households with slightly lower income move there. Ellen and Meltzer (2024) also find that lower-income households moved into the areas of New York that were hit harder by Hurricane Sandy.
@@
First, Section 2 describes the data from Intesa Sanpaolo, the data from ISPRA and Copernicus spatially matched to transacted homes, as well as EM-DAT dataset used to quantify our awareness measure. Next, Section 3 describes our econometric approach. The main results are in Section 4. Section 5 concludes.